A temperature of 10 degrees corresponds to a temperature of 50 degrees. Also 40 degrees corresponds to 104 degrees. Find the linear equation relating F and C.
Answer by nerdybill(7384)

A temperature of 10 degrees corresponds to a temperature of 50 degrees. Also 40 degrees corresponds to 104 degrees. Find the linear equation relating F and C.
.
You are given two points:
(10,50)
(40,104)
.
Slope:
(y2-y1)/(x2-x1)
= (104-50)/(40-10)
= 54/30
= 27/15
= 9/5
.
Use one of the two points say (10,50) and the slope (9/5) and plug it into the "point-slope" form:
y - y1 = m(x - x1)
y - 50 = (9/5)(x - 10)
y - 50 = (9/5)x - (9/5)10
y - 50 = (9/5)x - 18
y = (9/5)x + 32